# ORACLE-Grasp: Zero-Shot Task-Oriented Robotic Grasping using Large Multimodal Models

### v1

@@ IV Conclusions
This work presented O-Grasp, a hybrid grasping framework that integrates the semantic reasoning capabilities of LMMs with classical computer vision techniques to achieve robust, task-relevant grasp prediction on unknown objects. By introducing a structured dual-prompt interaction, using SCP for high level reasoning and GRP for spatial selection, we transform the inherently open-ended grasping task into a multiple-choice problem that LMMs can solve reliably. Through pre-cutting the image into candidate regions, aggregating results with IoU-based weighting and PCA, and applying depth-based clearance checks, O-Grasp effectively overcomes the spatial resolution limitations typical of LMMs. Additional mechanisms such as orientation refinement and early stopping contribute to both robustness and computational efficiency, enabling high success rates in both synthetic and real-world grasping experiments.
@@
Despite these promising results, several limitations remain. The inference cost associated with LMMs is still substantial, posing challenges for real-time, high-frequency control loops. Future improvements could involve model compression, distillation, or leveraging more efficient architectures. Moreover, O-Grasp currently operates from a single camera viewpoint at inference time. While the model can semantically reason about objects and select plausible grasp regions, it does not actively seek alternative perspectives when critical features are occluded. Future extensions incorporating multi-view fusion or active exploration strategies could significantly enhance the system’s adaptability, bringing robotic grasping closer to human-like flexibility.

### v2

@@ 5 Conclusions
This work presented ORACLE-Grasp, a hybrid grasping framework that integrates the semantic reasoning capabilities of LMMs with classical computer vision techniques to achieve robust, task-relevant grasp prediction on unknown objects. By introducing a structured dual-prompt interaction, using SCP for high-level reasoning and GRP for spatial selection, we transform the inherently open-ended grasping task into a multiple-choice problem that LMMs can solve reliably. Through pre-cutting the image into candidate regions, aggregating results with IoU-based weighting and PCA, and applying depth-based clearance checks, ORACLE-Grasp effectively overcomes the spatial resolution limitations typical of LMMs. Additional mechanisms such as orientation refinement and early stopping contribute to both robustness and computational efficiency, enabling high success rates in both synthetic and real-world grasping experiments.
While traditional state-of-the-art grasp synthesis methods demonstrate high precision on known object distributions, they differ fundamentally from ORACLE-Grasp in both objective and infrastructure. Supervised models are typically ’expert students’ of specific datasets; in contrast, ORACLE-Grasp acts as a ’reasoning generalist’ that requires no prior training on specific object classes. Furthermore, standard geometric methods often prioritize mechanical stability over functional suitability. For instance, a purely geometric approach might successfully grasp a drill by its bit or a screwdriver by its shaft, which precludes subsequent task execution. ORACLE-Grasp prioritizes affordance alignment, using LMM-driven reasoning to identify handles and ergonomic interfaces. Finally, whereas many SOTA frameworks rely on dense 3D volumetric reconstruction or complex point cloud processing, our framework performs the core localization and orientation reasoning in the RGB domain. Depth information is utilized strictly for metric execution and local clearance checks. This shift from 3D-heavy modeling to LMM-driven visual reasoning allows ORACLE-Grasp to operate with lower computational overhead and greater flexibility in unstructured, zero-shot environments.
@@
Despite these promising results, several limitations remain. The inference cost associated with LMMs is still substantial, posing challenges for real-time, high-frequency control loops. Future improvements could involve model compression, distillation, or leveraging more efficient architectures. Moreover, ORACLE-Grasp currently operates from a single camera viewpoint at inference time. While the model can semantically reason about objects and select plausible grasp regions, it does not actively seek alternative perspectives when critical features are occluded. Future extensions incorporating multi-view fusion or active exploration strategies could significantly enhance the system’s adaptability, bringing robotic grasping closer to human-like flexibility.
While the current validation of ORACLE-Grasp focuses on single-object scenes to isolate its zero-shot affordance reasoning, the framework’s modular design is inherently extensible in future work to cluttered, multi-object environments. By parameterizing the SCP to accept specific target labels or complex human instructions, ranging from explicit commands like ”pick-up the blue wrench” to implicit, task-oriented goals like ”I need to tighten this bolt”, the system can transition from autonomous principal object identification to targeted semantic search. In these scenarios, the LMM acts as a high-level planner that identifies the appropriate tool and its functional parts, allowing the subsequent GRP to ignore surrounding clutter and focus on local affordances. This evolution effectively bridges the gap between abstract natural language intent and low-level physical execution, paving the way for intuitive service robots capable of reasoning through the partial occlusions and complex spatial relationships of real-world settings.
[…]
Supplementary information
Videos demonstrating the robot experiments, both in simulation and real-world environments, are available in the supplementary material.
Declarations
[…]
Funding
This work was supported by the Israel Science Foundation (No. 451/24).
[…]
Conflict of interest The authors declare that there is no conflict of interest or Competing interests.
[…]
Data availability The test data collected in this work will be made available on request.
[…]
Code availability Code will be made available on request.